Find three consecutive odd integers such that 4 times the first is 8 less than 3 times the sum of the last two.
Answer by edjones(8007)

x, x+2, x+4 are the integers
4x+8=3(x+2+x+4)
...=3(2x+6)
...=6x+18
-2x=10
x=-5
-5, -3 and -1 are the integers
.
Ed